# Improving Sensitivity of Vector-like Top Partner Searches with Jet Substructure

### v1

@@ II.3 Background processes
Several SM processes that constitute the same final state as that of the signal are considered as backgrounds. The background MC samples are obtained from the CMS Open Data. Table 2 summarizes these background processes together with their respective cross sections. In the case of $\mathrm{t\bar{t}}$ + jets background, MC samples binned in $\text{H}_{\text{T}}$ are stitched with the so-called inclusive sample according to the method described in Ref. Ehatäht and Veelken (2022) for better statistical modelling of the boosted topology. Table 3 summarizes the stitching weights applied to events lying in different $\text{H}_{\text{T}}$ bins. For the QCD multijet background, a combination of different MC samples in non-overlapping $\hat{\text{p}}_{\text{T}}$ bins above 151515~{}15GeV is considered.
Table 2: List of SM backgrounds along with their respective cross sections and MC sample sizes.
@@
[pb] (in Millions)
$\mathrm{t\bar{t}}$ + jets ttx (2021) 833.900 Czakon and Mitov (2014) NNLO + NNLL 117.70
t($\bar{\text{t}}$)W + jets tWj (2021) 79.300 Kidonakis and Yamanaka (2021) NNLO 2.00
$\mathrm{t\bar{t}Z}$ttZ (2021) 0.863 Kulesza et al. (2019) NLO + NNLL 22.16
$\mathrm{t\bar{t}W}$ttW (2021) 0.745 Buonocore et al. (2023) NNLO (QCD) + NLO (EW) 0.83
WZ + jets WZj (2021) 56.850 Grazzini et al. (2016a) NNLO 13.19
WW + jets WWj (2021) 1.371 Grazzini et al. (2016b) NNLO 8.92
QCD multijet QCD (2021) $1.981\times 10^{9}$ LO 97.19
Table 3: Stitching weights for $\mathrm{t\bar{t}}$ + jets events in different $\text{H}_{\text{T}}$ bins for 300 fb-1 integrated luminosity.
@@
×$\times 10^{-2}$$\times 10^{-2}$$\times 10^{-3}$$\times 10^{-4}$
@@ III.1 Dynamic radius jet clustering algorithm
The jet clustering algorithms are crucial steps in the analysis of events in a hadron collider experiment such as the LHC Salam (2010); Sterman and Weinberg (1977). The currently used algorithms are fixed-radius sequential recombination inclusive jet algorithms, namely the $\text{k}_{\text{T}}$ (KT), anti – $\text{k}_{\text{T}}$ (AK), and Cambridge – Aachen (CA) Ellis and Soper (1993); Cacciari et al. (2008); Wobisch and Wengler (1998) algorithms. These traditional algorithms take a list of four-momenta of the particles produced in an event by the process of recombining one or more of these particles into several jets. In their iterative steps, these algorithms calculate pair-wise “$\text{k}_{\text{T}}$” distance ($\text{d}_{\text{ij}}$) between $\text{i}^{\text{th}}$ and $\text{j}^{\text{th}}$ particles and $\text{i}^{\text{th}}$ particle “beam” distance ($\text{d}_{\text{iB}}$), defined as
$\text{d}_{\text{ij}}=\min\left(\text{p}_{\text{T,\ i}}^{2n}\ ,\ \text{p}_{\text{T,\ j}}^{2n}\right)\ \Delta\text{R}_{\text{ij}}^{2},\ \text{and}$ (10)
$\text{d}_{\text{iB}}=\text{p}_{\text{T,\ i}}^{2n}\ \text{R}^{2}$ (11)
[…]
@@
In this study, we use the “Dynamic radius” jet clustering algorithm Mukhopadhyaya et al. (2023a), which attempts to dynamically modify the jet radius during the evolution of each jet through the iterative steps of the algorithm. The approach is to start from an initial radius $R_{0}$, the radius for each evolving “proto-jet” gets modified depending on its internal structure during the process of clustering. The square of the additive modifier $\sigma_{i}$ that is added to each evolving proto-jet at each interactive stage is given by
$\sigma_{i}^{2}=\dfrac{\displaystyle\sum_{a<b}\text{p}_{\text{T}}^{a}\,\text{p}_{\text{T}}^{b}\ \Delta\mathrm{R}_{ab}^{2}}{\displaystyle\sum_{a<b}\text{p}_{\text{T}}^{a}\,\text{p}_{\text{T}}^{b}}-\left(\dfrac{\displaystyle\sum_{a<b}\text{p}_{\text{T}}^{a}\,\text{p}_{\text{T}}^{b}\ \Delta\mathrm{R}_{ab}}{\displaystyle\sum_{a<b}\text{p}_{\text{T}}^{a}\,\text{p}_{\text{T}}^{b}}\right)^{2},$ (12)
@@
where $a$ and $b$ run over the constituents of the pseudojet. So, the final radius for the $i^{\rm th}$ jet becomes $R_{\rm fin}=R_{0}+\sigma_{i}$. The advantage one gets in this method is that the radius is no longer fixed; it gets adapted to the requirement of the jet that is being formed. The modifier is a slight variation of a combination of a few jet substructure observable ECFs Larkoski et al. (2013), which tends to become larger for top or heavy boson jets Aaboud et al. (2019). On the other hand, it tends to be smaller for narrow jets like the ones appearing from light quarks or gluons. Since our study involves one fat jet along with narrow $b$-jets, the efficiency of the dynamic radius jet clustering is expected to be better.
@@ III.2 Jet substructure and parameter choices
  • Soft Drop Grooming: To mitigate contamination from pileup, underlying events, etc., inside a jet, and improve the reliability of the jet substructure measurements, we have used the Soft Drop de-clustering method Larkoski et al. (2014a). This method primarily aims to remove soft and wide-angle radiations inside a jet while retaining the hard core of the jet. To briefly describe, the Soft Drop algorithm is applied on a jet that has been re-clustered using a sequential clustering algorithm such as CA, KT, or AK. The essential step of this method is to remove soft and wide-angle subjet after de-clustering the jet into two subjets (labelled as $i$ and $j$) by checking the following condition.
    $\frac{\min(p_{T,i},p_{T,j})}{p_{T,i}+p_{T,j}}<z^{\rm SD}_{\text{cut}}\left(\frac{\Delta R_{ij}}{R^{\rm SD}}\right)^{\beta^{\rm SD}}$ (13)
    Here, $p_{T,i}$ and $p_{T,j}$ are the transverse momenta of the subjets, $\Delta R_{ij}$ is the distance between them in the $\eta$-$\phi$ plane. This process is repeated until the condition fails to be satisfied. The radius $R^{\rm SD}$ is the characteristic jet radius usually taken to be the radius parameter used during the clustering. In case of the dynamic radius jet clustering algorithm, we have taken this parameter to be $R_{d}=R_{0}+\sigma$, the final dynamically adjusted radius for each jet.
@@
    The two parameters of Soft Drop are (1) $z^{\rm SD}_{\text{cut}}$: energy fraction threshold and (2) $\beta^{\rm SD}$: angular exponent regulating the angular separation between the two subjets. In the special case of $\beta=0$, the method reduces to the Modified Mass-Drop Tagger Butterworth et al. (2008); Dasgupta et al. (2013). The grooming then depends only on the momentum fraction $z$ and not on the angular separation. We have used the value of $\beta^{\rm SD}=0.0$ to keep our analysis aligned with the available jet substructure measurement Aaboud et al. (2019). We also check the alternative value of $1.0$. For the energy fraction threshold parameter, a nominal value of $z_{\rm cut}^{\rm SD}=0.1$ is used to match the studies performed in Refs. Larkoski et al. (2014a); Larkoski and Thaler (2014); Larkoski et al. (2015); Sirunyan et al. (2018a).
@@
  • $N$-subjettiness: The jet substructure method $N$-subjettiness Thaler and Van Tilburg (2011), provides a measure of the consistency of jet having $N$ subjets inside. In our event topology, we expect boosted top fat jet, which gives rise to a three-prong structure corresponding to its decay products: $t\to bW\to bq\bar{q}$. Therefore, this method becomes particularly useful with the value of $N=3$. The $N$-subjettiness observable captures these subjets by fitting $N$ subjet axes inside the fat jet. This is performed by minimizing the $N$-subjettiness measure Thaler and Van Tilburg (2012) defined as
@@
    $\tilde{\tau}_{N}^{(\beta^{\rm NS})}=\frac{1}{d_{0}}\sum_{i\in{\rm jet}}{p_{T}}_{i}\min\left\{\Delta R_{i,a_{1}}^{\beta^{\rm NS}},\Delta R_{i,a_{2}}^{\beta^{\rm NS}},\cdots,\Delta R_{i,a_{N}}^{\beta^{\rm NS}}\right\},$ (14)
@@
    where the sum runs over all constituents in the jet, and $\Delta R_{i,a_{k}}$ is the distance in $\eta$-$\phi$ plane between constituent $i$ and subjet axis $a_{k}$. The prefactor $1/d_{0}$ is a normalisation factor. The exponent $\beta^{\rm NS}$ and the method used to determine the seed $N$ axes (e.g., exclusive $k_{T}$ axes from an iterative algorithm) are parameters of choice. We have used $N$-subjettiness method implemented in Fastjet Contrib Cacciari et al. (2012). In our study, we use the ratios $\tau_{32}=\tau_{3}/\tau_{2}$ and $\tau_{21}=\tau_{2}/\tau_{1}$ as a top tagger with $\beta^{\rm NS}=1.0$ and the seed axes obtained via WTA_KT_Axes method Bertolini et al. (2014); Larkoski et al. (2014b); Salam .
@@
  • Energy Correlation Functions: These are a class of infrared and collinear safe observables proposed to study jet properties Larkoski et al. (2013). The first three lowest point functions take the form
    ${\rm ECF}_{1}(\beta^{\rm ECF})$ $=$ $\sum_{i\in J}{p_{T}}_{i}\ ,$ (15)
    ${\rm ECF}_{2}(\beta^{\rm ECF})$ $=$ $\sum_{i<j\in J}{p_{T}}_{i}\,{p_{T}}_{j}\Delta R_{ij}^{\beta^{\rm ECF}}\ ,$ (16)
    ${\rm ECF}_{3}(\beta^{\rm ECF})$ $=$ $\sum_{i<j<k\in J}{p_{T}}_{i}{p_{T}}_{j}{p_{T}}_{k}\left(\Delta R_{ij}\Delta R_{ik}\Delta R_{jk}\right)^{\beta^{\rm ECF}}\ ,$ (17)
@@
    where the sum is over the jet constituents and the exponent $\beta^{\rm ECF}$ is a parameter that gives the weight to the angular separation between the jet constituents. These functions help distinguish between jets originating from different partonic configurations. In case of top quark study, a better discrimination from QCD jets is obtained for $\beta^{\rm ECF}=1.0$Larkoski et al. (2013). We therefore choose to use $\beta^{\rm ECF}=1.0$ as a nominal value and $\beta^{\rm ECF}=0.5$ for an alternate value.
    Instead of using the ECFs as defined in Eqs.(15) – (17) directly, their ratios are better suited as they remove the $p_{T}$ dependence. The following ratios Aaboud et al. (2019) are studied.
@@
    $\quad C_{2}=\displaystyle\frac{e_{3}}{(e_{2})^{2}}\ ,\qquad\text{and}\ \qquad$ $D_{2}=\frac{e_{3}}{(e_{2})^{3}}\ .$ (19)
The dynamic radius jet algorithm, along with the jet substructure methods, has been implemented within Delphes and CMS SoftWare version 7_6_7 as appropriate plugins. The choices of parameters for the jet substructure methods described above for the dynamic radius jets, including both nominal and alternate values used for variations, are summarized in Table 4. The impact of these choices is studied by comparing the shapes of different observables sensitive to the respective substructure methods, as shown in Fig. 4. As seen in the figure, the distributions for nominal and alternate parameter choices look quite similar, suggesting that the overall analysis remains fairly robust against these variations. In addition, the nominal choice for the dynamic radius jet algorithm is also compared with the fixed-radius counterpart, as shown in Fig. 5.
Table 4: Choice of various parameters for dynamic radius fat jets.
@@
Value Value
Soft Drop Larkoski et al. (2014a) $\beta^{\text{SD}}$ 0.0 Aaboud et al. (2019) 1.0
$z_{\text{cut}}^{\text{SD}}$ 0.1 Aaboud et al. (2019)
$N$-subjettiness Thaler and Van Tilburg (2011) $\beta^{\text{NS}}$ 1.0 Aaboud et al. (2019)
Seed Axes Stewart et al. (2015) WTA_KT_Axes Larkoski and Thaler (2014)
ECF Larkoski et al. (2013) $\beta^{\text{ECF}}$ 1.0 Aaboud et al. (2019) 0.5
Figure 4: Comparison of Soft Drop mass (upper-left), $\tau_{21}$ (upper-middle), $\tau_{32}$ (upper-right), and the ECF ratios $C_{2}$ (lower-left), $D_{2}$ (lower-middle), and $e_{3}$ (lower-right) distributions of the leading fat jet corresponding to nominal and alternate values of substructure parameters for dynamic radius jets in $\mathrm{t\bar{t}}$ + jets events.
Figure 5: Comparison of Soft Drop mass (left), $\tau_{21}$ (middle), and $\tau_{32}$ (right) distributions of the leading fat jet between fixed- and dynamic radius algorithms in $\mathrm{t\bar{t}}$ + jets events.

### v2

@@ II.3 Background Processes
[…]
accuracy [pb] accuracy [Million]
$\mathrm{t\bar{t}}$ + jets ttxj_full NLO Alwall:2014hca ; Frederix:2012ps ; Sjostrand:2007gs 833.900 Czakon:2011xx NNLO + NNLL 117.70
[…]
$\mathrm{t\bar{t}Z}$ ttZjets_full LO Alwall:2014hca ; MLM01 ; Mangano:2006rw ; Sjostrand:2007gs 0.863 Kulesza:2018tqz NLO + NNLL 22.16
$\mathrm{t\bar{t}W}$ ttWjets_full NLO Alwall:2014hca ; Frederix:2012ps ; Sjostrand:2007gs 0.745 Buonocore:2023ljm NNLO (QCD) 0.83
+ NLO (EW)
WZ + jets WZjets_full NLO Alwall:2014hca ; Frederix:2012ps ; Sjostrand:2007gs 56.850 Grazzini:2016swo NNLO 13.19
WW + jets WWjets_full NLO Nason:2013ydw ; Alioli:2010xd ; Sjostrand:2007gs 1.371 Grazzini:2016ctr NNLO 8.92
QCD multijet QCD_full LO Sjostrand:2007gs $1.981\times 10^{9}$ LO 97.19
Table 4: Stitching weights for $\mathrm{t\bar{t}}$ + jets events in different $\text{H}_{\text{T}}$ bins for 300 fb-1 integrated luminosity, computed following the prescription in Ref. Ehataht:2021rkh .
@@
$\times 10^{-2}$ $\times 10^{-2}$ $\times 10^{-3}$ $\times 10^{-4}$
@@ III.1 Dynamic Radius Jet Clustering Algorithm
The jet clustering algorithms are crucial steps in the analysis of events in a hadron collider experiment such as the LHC Salam:2010nqg ; Sapeta:2015gee . The currently used algorithms are fixed-radius sequential recombination inclusive jet algorithms, namely the $\text{k}_{\text{T}}$ (KT), anti-$\text{k}_{\text{T}}$ (AK), and Cambridge-Aachen (CA) Ellis:1993tq ; Catani:1993hr ; Cacciari:2008gp ; Dokshitzer:1997in ; Wobisch:1998wt algorithms. These traditional algorithms take as input the list of four momenta of the particles from an event and successively recombine them into several jets. In their iterative steps, these algorithms calculate pair-wise $\text{k}_{\text{T}}$” distance ($\text{d}_{\text{ij}}$) between $\text{i}^{\text{th}}$ and $\text{j}^{\text{th}}$ four momenta and “beam” distance ($\text{d}_{\text{iB}}$) for the $\text{i}^{\text{th}}$ four momentum, defined as
$\text{d}_{\text{ij}}=\min\left(\text{p}_{\text{T,\ i}}^{2n}\ ,\ \text{p}_{\text{T,\ j}}^{2n}\right)\ \frac{\Delta\text{R}_{\text{ij}}^{2}}{\text{R}^{2}},\ \text{and}$ (10)
$\text{d}_{\text{iB}}=\text{p}_{\text{T,\ i}}^{2n}$ (11)
to decide whether to merge two four momenta or declare a particular four momentum as a final jet. Here, $\Delta\text{R}_{\text{ij}}$ is the distance between $\text{i}^{\text{th}}$ and $\text{j}^{\text{th}}$ four momenta in the pseudorapidity-azimuth ($\eta$-$\phi$) plane. The two four momenta are merged to give a new four momentum if $\text{d}_{\text{ij}}$ becomes the least among all the pair-wise $\text{k}_{\text{T}}$” and beam distances. Otherwise, the $\text{i}^{\text{th}}$ four momentum corresponding to the least beam distance is declared as a final jet. Importantly, the algorithm has parameters $n$, which usually takes values 1, 0,and 1-1,\ 0,\ $\text${and}\ 1 corresponding to AK, CA, and KT algorithms respectively, and R as the distance measure in the $\eta$-$\phi$ plane. These algorithms typically output jets of a fixed size, which is determined by the input parameter R to the algorithm. In this article, the AK algorithm is chosen as the default since AK jets have regular conical shape with an well-defined jet area Cacciari:2008gp .
@@
The traditional jet clustering approaches mentioned above have the limitation in producing varying-sized jets in an event for a given value of R. Several interesting attempts have been made to modify the algorithms to accommodate jets of variable sizes; such as “Variable R” jets Krohn:2009zg , “Fuzzy” jets Mackey:2015hwa , “SHAPER” Ba:2023hix , “XCone” Stewart:2015waa , “PAIReD” jets Mondal:2023law , “dynamic radius” jets Mukhopadhyaya:2023rsb etc. Here, we focus on the “dynamic radius” jet clustering algorithm Mukhopadhyaya:2023rsb , which attempts to dynamically modify the jet radius during the evolution of each jet through the iterative steps of the algorithm. The approach is to start from an initial radius $\rm R_{0}$}, the radius for each evolving “proto-jet” gets modified depending on its internal structure during the process of clustering. The square of the additive modifier $\sigma_{i}$ to $\rm R_{0}$} for each evolving “proto-jet” at each interactive stage is given by
$\sigma_{\rm i}^{2}=\dfrac{\displaystyle\sum_{a<b}{\rm p}_{{\rm T},a}\,{\rm p}_{{\rm T},b}\ \Delta\mathrm{R}_{ab}^{2}}{\displaystyle\sum_{a<b}{\rm p}_{{\rm T},a}\,{\rm p}_{{\rm T},b}}-\left(\dfrac{\displaystyle\sum_{a<b}{\rm p}_{{\rm T},a}\,{\rm p}_{{\rm T},b}\ \Delta\mathrm{R}_{ab}}{\displaystyle\sum_{a<b}{\rm p}_{{\rm T},a}\,{\rm p}_{{\rm T},b}}\right)^{2},$ (12)
@@
where $a$ and $b$ run over the constituents of the proto-jet. So, the final radius for the $i^{\rm th}$ jet becomes $\mathrm{R_{d}}=\mathrm{R_{0}+\sigma_{i}}$. The advantage one gets in this method is that the radius is no longer fixed; it gets adapted to the requirement of the jet that is being formed. The modifier is a slight variation of a combination of a few jet substructure observable ECFs Larkoski:2013eya , which tends to become larger for top or heavy boson jets ATLAS:2019kwg . On the other hand, it tends to be smaller for narrow jets like the ones appearing from light quarks or gluons. Since our study involves one fat jet along with narrow b-jets, the efficiency of the DR jet clustering is expected to be better.
@@ III.2 Jet Substructure and Parameter Choices
To mitigate contamination from pile up, underlying events etc., inside a jet, and improve the reliability of the jet substructure measurements, we have used the Soft Drop de-clustering method Larkoski:2014wba . This method primarily aims to remove soft and wide-angle radiations inside a jet while retaining the hard core of the jet. To briefly describe, the Soft Drop algorithm is applied on a jet that has been re-clustered using the CA algorithm (which orders by angle). The essential step of this method is to remove soft and wide-angle subjet after de-clustering the jet into two subjets (labelled as i and j) by checking the following condition.
$\mathrm{\frac{\min(p_{T,i},p_{T,j})}{p_{T,i}+p_{T,j}}}<z^{\rm SD}_{\text{cut}}\left(\mathrm{\frac{\Delta R_{ij}}{{\rm R}^{\rm SD}}}\right)^{\beta^{\rm SD}}$ (13)
@@
Here, $\mathrm{p_{T,i}}$ and $\mathrm{p_{T,j}}$ are the transverse momenta of the subjets, $\mathrm{\Delta R_{ij}}$ is the distance between them in the $\eta$-$\phi$ plane. This process is repeated until the condition fails to be satisfied. The radius ${\rm R}^{\rm SD}$ is the characteristic jet radius usually taken to be the radius parameter used during the clustering. In case of the DR jet clustering algorithm, we have taken this parameter to be $\mathrm{R_{d}=R_{0}+\sigma}$, the final dynamically adjusted radius for each jet. The two parameters of Soft Drop are (1) $z^{\rm SD}_{\text{cut}}$: the energy fraction threshold and (2) $\beta^{\rm SD}$: the exponent regulating the angular separation between the two subjets. In the special case of $\beta^{\rm SD}=0$, the method reduces to the Modified Mass-Drop Tagger Butterworth:2008iy ; Dasgupta:2013ihk . The grooming then depends only on the momentum fraction $z$ and not on the angular separation. We have used the value of $\beta^{\rm SD}=0$ to keep our analysis aligned with the available jet substructure measurement ATLAS:2019kwg . We also check the alternative value of $\beta^{\rm SD}=1$. For the energy fraction threshold parameter, a nominal value of $z_{\rm cut}^{\rm SD}=0.1$ is used to match the studies performed in Refs. Larkoski:2014wba ; Larkoski:2014bia ; Larkoski:2015lea ; CMS:2017qlm .
NN-subjettiness
@@
The jet substructure method $N$-subjettiness Thaler:2010tr , provides a measure of the consistency of a jet having $N$ subjets inside. In our event topology, we expect boosted top, W or Z fat jet, which gives rise to a three- or two-prong structure. Therefore, this method becomes particularly useful with the value of $N=3$ and 2. The $N$-subjettiness observable captures these subjets by fitting $N$ subjet axes inside the fat jet. This is performed by minimizing the $N$-subjettiness measure Thaler:2011gf defined as
@@
$\mathrm{\tilde{\tau}_{N}^{(\beta^{\rm NS})}=\frac{1}{d_{0}}\sum_{i\in{\rm jet}}{p_{T}}_{i}}\min\left\{\Delta{\rm R}_{{\rm i},a_{1}}^{\beta^{\rm NS}},\Delta{\rm R}_{{\rm i},a_{2}}^{\beta^{\rm NS}},\cdots,\Delta{\rm R}_{{\rm i},a_{N}}^{\beta^{\rm NS}}\right\},$ (14)
@@
where the sum runs over all constituents in the jet, and $\Delta{\rm R}_{{\rm i},a_{k}}$ is the distance in $\eta$-$\phi$ plane between constituent i and subjet axis $a_{k}$. The prefactor $\rm 1/d_{0}$} is a normalization factor. The minimization is performed over the choices of axes, and the final minimum value of $\tilde{\tau}_{N}$, denoted as $\tau_{N}$, represents the optimal measure of $N$-subjettiness. The exponent $\beta^{\rm NS}$ and the method used to determine the seed $N$ axes (e.g., exclusive $\text{k}_{\text{T}}$ axes from an iterative algorithm) are parameters of choice. We have used $N$-subjettiness method implemented in Fastjet Contrib Cacciari:2011ma . In our study, we use $\beta^{\rm NS}=1$, the seed axes obtained by WTA_KT_Axes method Bertolini:2013iqa ; Larkoski:2014uqa ; Salam_unpub , and the ratios $\tau_{32}=\tau_{3}/\tau_{2}$ and $\tau_{21}=\tau_{2}/\tau_{1}$ to respectively identify three- and two-prong structures inside a fat jet.
Energy Correlation Functions
@@
These are a class of infrared and collinear safe observables proposed to study jet properties Larkoski:2013eya . The first three lowest point functions take the form
${\rm ECF}_{1}(\beta^{\rm ECF})$ $=$ $\mathrm{\sum_{i\,\in\,jet}p_{T,i}}\ ,$ (15)
${\rm ECF}_{2}(\beta^{\rm ECF})$ $=$ $\mathrm{\sum_{i<j\,\in\,jet}p_{T,i}\,p_{T,j}\,\Delta R_{ij}^{\beta^{\rm ECF}}}\ ,$ (16)
${\rm ECF}_{3}(\beta^{\rm ECF})$ $=$ $\mathrm{\sum_{i<j<k\,\in\,jet}p_{T,i}\,p_{T,j}\,p_{T,k}\left(\Delta R_{ij}\,\Delta R_{ik}\,\Delta R_{jk}\right)^{\beta^{\rm ECF}}}\ ,$ (17)
@@
where the sum is over the jet constituents and the exponent $\beta^{\rm ECF}$ is a parameter that gives the weight to the angular separation between the jet constituents. These functions help distinguish between jets originating from different partonic configurations. In case of top quark study, a better discrimination from QCD jets is obtained for $\beta^{\rm ECF}=1.0$Larkoski:2013eya . We therefore choose to use $\beta^{\rm ECF}=1.0$ as a nominal value and $\beta^{\rm ECF}=0.5$ for an alternate value. Instead of using the ECFs as defined in Eqs. (15)–(17) directly, their ratios are better suited as they remove the $\text{p}_{\text{T}}$ dependence. The following ratios ATLAS:2019kwg are studied.
@@
$\quad C_{2}=\displaystyle\frac{e_{3}}{(e_{2})^{2}}\ ,\qquad\text{and}\ \qquad\$$D_{2}=\frac{e_{3}}{(e_{2})^{3}}\ .$ (19)
The DR jet algorithm, along with the jet substructure methods, has been implemented within Delphes and CMS SoftWare version 7_6_7 as appropriate plugins. The choices of parameters for the jet substructure methods described above for the DR jets, including both nominal and alternate values used for variations, are summarized in Table 5. The impact of these choices is studied by comparing the shapes of different observables sensitive to the respective substructure methods, as shown in Figure 4, both at truth and reconstructed levels. As seen in the figure, the distributions for nominal and alternate parameter choices look quite similar, suggesting that the overall analysis remains fairly robust against these variations. In addition, substructure observables of the DR jets are also compared with the fixed-radius counterpart at the reconstructed level, and are shown in Figure 5. Only the nominal values of the substructure parameters are considered for this comparison. Figure 5 demonstrates that the distributions of $\tau_{21}$ and $\tau_{32}$ for the DR clustering are more shifted towards 0 compared to the fixed-radius clustering, indicating better reconstruction of the hadronically decaying boosted W bosons and top quarks in the former case with a fat jet Thaler:2010tr .
Table 5: Parameter choices for jet substructure methods applied to the DR fat jets. “Nominal Values” correspond to those used in the final analysis, while “Alternate Values” are considered for the control study shown in Figure 4.
@@
Values Values
Soft Drop Larkoski:2014wba $\beta^{\text{SD}}$ 0.0 ATLAS:2019kwg 1.0
$z_{\text{cut}}^{\text{SD}}$ 0.1 ATLAS:2019kwg 0.1
NN-subjettiness Thaler:2010tr $\beta^{\text{NS}}$ 1.0 ATLAS:2019kwg 1.0
  Seed Axes Stewart:2015waa   WTA_KT_Axes Larkoski:2014bia   WTA_KT_Axes
ECF Larkoski:2013eya $\beta^{\text{ECF}}$ 1.0 ATLAS:2019kwg 0.5
Figure 4: Comparisons of Soft Drop mass (upper-left), $\tau_{21}$ (upper-right), $\tau_{32}$ (middle-left), and the ECF ratios $C_{2}$ (middle-right), $D_{2}$ (lower-left), and $e_{3}$ (lower-right) distributions of the leading anti-$\text{k}_{\text{T}}$ fat jet corresponding to nominal and alternate values of substructure parameters as indicated in Table 5 for dynamic radius jet clustering algorithm at truth and reconstructed levels in $\mathrm{t\bar{t}}$ + jets events.
Figure 5: Comparisons of Soft Drop mass (upper-left), $\tau_{21}$ (upper-right), and $\tau_{32}$ (lower) distributions of the reconstructed leading anti-$\text{k}_{\text{T}}$ fat jet between fixed-radius (${\rm R}=0.8$) and dynamic radius jet clustering algorithms in $\mathrm{t\bar{t}}$ + jets events are presented. The nominal choice of substructure parameters as indicated in Table 5 has been used for the dynamic radius algorithm to obtain these plots.